OpenStudy (anonymous):

If f(x)=10x - 10, what is the value of (f °f^-1)(-10)?

OpenStudy (isaidavila):

(F°F^-1)=10(1/10x-10)-10 I think culd be that

OpenStudy (michele_laino):

hint: we have this: \[\Large f\left( 0 \right) = - 10 \Rightarrow {f^{ - 1}}\left( { - 10} \right) = 0\]

OpenStudy (anonymous):

So f^-1 is 10? Sorry, I don't really get this

OpenStudy (michele_laino):

by definition, we have: \[\Large \left( {f \circ {f^{ - 1}}} \right)\left( { - 10} \right) = f\left\{ {{f^{ - 1}}\left( { - 10} \right)} \right\} = ...\]

OpenStudy (michele_laino):

another step: \[\Large \left( {f \circ {f^{ - 1}}} \right)\left( { - 10} \right) = f\left\{ {{f^{ - 1}}\left( { - 10} \right)} \right\} = f\left( 0 \right) = ...?\]

OpenStudy (anonymous):

Is it 0? Because anything times 0 is 0?

OpenStudy (dumbcow):

a function and its inverse cancel themselves out \[f(f^{-1} (x)) = x\] \[f^{-1}(f(x)) = x\]

OpenStudy (anonymous):

I'll keep that in mind. Thanks 🐢 I just put 0, since it wasn't for a grade anyways lol The answer was apparently -10. Could you please explain why?